# FlashBias: Fast Computation of Attention with Bias

### v1

@@ Abstract
Attention mechanism has emerged as a foundation module of modern deep learning models and has also empowered many milestones in various domains. Moreover, FlashAttention with IO-aware speedup resolves the efficiency issue of standard attention, further promoting its practicality. Beyond canonical attention, attention with bias also widely exists, such as relative position bias in vision and language models and pair representation bias in AlphaFold. In these works, prior knowledge is introduced as an additive bias term of attention weights to guide the learning process, which has been proven essential for model performance. Surprisingly, despite the common usage of attention with bias, its targeted efficiency optimization is still absent, which seriously hinders its wide applications in complex tasks. Diving into the computation of FlashAttention, we prove that its optimal efficiency is determined by the rank of the attention weight matrix. Inspired by this theoretical result, this paper presents FlashBias based on the low-rank compressed sensing theory, which can provide fast-exact computation for many widely used attention biases and a fast-accurate approximation for biases in general formalization. FlashBias can fully take advantage of the extremely optimized matrix multiplication operation in modern GPUs, achieving 1.5$\times$ speedup for AlphaFold, and over 2$\times$ speedup for attention with bias in vision and language models without loss of accuracy.
@@ 1 Introduction
In FlashAttention dao2022flashattention , researchers find that instead of computation FLOPS, the read-write (IO) overload of GPU high bandwidth memory (HBM) is the actual bottleneck of speed. As bias matrices are usually dense, it is really hard to bypass the quadratic IO complexity, where the computation needs to read the whole bias term from HBM at least once, making its speedup intractable. In this paper, we notice that the IO challenge that we face here can be recast into the classical compressed sensing problem donoho2006compressed , whose basic assumption is that the “measurement” (corresponding to IO overload here) is expensive but the computation (corresponding to the fast on-chip computation) is cheap. This new perspective offers us valuable theoretical understandings. Specifically, for a dense matrix, the optimal “measurement” (IO overload) is highly related to the matrix’s rank candes2010power . This inspires us to dive into the rank of attention bias and surprisingly find that most of the widely used biases are inherently of low rank, eventually discovering one flexible way to enable fast computation of attention with bias.
@@
  • FlashBias can accelerate a family of widely-used backbones without loss of accuracy, which brings 1.5$\times$ speedup for AlphaFold and over 2$\times$ speedup for vision and language models.
@@ 3.1 Rethinking FlashAttention Computation
Given $\mathbf{q}\in\mathbb{R}^{N\times C},\mathbf{k},\mathbf{v}\in\mathbb{R}^{M\times C}$, bias $\mathbf{b}$ of rank $R$ and SRAM of size $S$ where $(C+R)\leq S\leq N(C+R)$ , there does not exist an algorithm to compute exact attention with bias through $o\big{(}\frac{NM(C^{2}+R^{2})}{S}\big{)}$ HBM access for all $S\in[(C+R),N(C+R)]$ Here $o(\ast)$ represents the strict asymptotic upper bound.
@@ 3.2 FlashBias
where [|][$\ast$|$\ast$][ ∗ | ∗ ] denotes the concatenation operation along the channel dimension. Notably, this design significantly reduces the storage cost for the attention bias from $\mathcal{O}(NM)$ to $\mathcal{O}\left((N+M)R\right)$. Although its design will require recalculating the bias weight, this computation is just a simple matrix multiplication of $\phi_{\mathbf{q}}(\mathbf{x}_{\mathbf{q}})\phi_{\mathbf{k}}(\mathbf{x}_{\mathbf{k}})^{\top}$, an operation that has been extremely optimized on modern GPUs.
@@
Some models such as Swin Transformer liu2021Swin and Pangu-Weather bi2023accurate adopt the learnable model parameters for relative position encoding. Specifically, each bias term in their model is a $N\times M$ matrix of model parameters. As this type of bias is fixed once the model has been well trained, it is convenient to conduct Singular Value Decomposition (SVD) klema1980singular for low-rank decomposition of these parameters. In practice, we precompute SVD once offline, incurring negligible runtime overhead. The resulting decomposed factor tensors can then be utilized to accelerate the subsequent inference process, thanks to their low-rank nature.
@@
Although SVD and neural decomposition will introduce approximation error, owing to the native low-rank property of attention bias in modern Transformers, it will not affect the final performance by setting $R$ as a reasonable value. For example, in some layers in Swin Transformer, $R=32$ can maintain over 99% energy of the original bias with shape $576\times 576$. More case studies are included in Appendix C and D.
@@
For FlashAttention with bias, its HBM access is $\Theta(\frac{NMC^{2}}{S}+NM)$ dao2023flashattention . Given a regular setting of Transformers and GPU ($C=64$, $S$ is 100KB) and supposing $R=64$, $N,M\gg C,R$, then HBM IO of FlashBias is $\approx 6$, smaller than FlashAttention.
@@ 4 Experiments
As summarized in Table 2, we extensively test FlashBias on Transformers for language modeling, computer vision and scientific problems, which can significantly speed up model efficiency at both training and test phases. This section will first present an overall efficiency comparison among different techniques and then integrate FlashBias to different domain-specific Transformers with various bias types. All the experiments were performed in PyTorch 2.5.0 Paszke2019PyTorchAI on a single A100 40GB GPU.
@@ 4.1 Overall Comparison
To give a clear comparison among FlashBias, vanilla FlashAttention with Bias dao2023flashattention and the latest FlexAttention dong2024flex , we make a comprehensive efficiency evaluation based on a plain Transformer NIPS2017_3f5ee243 , which consists of 8 layers. Each Transformer layer involves a feedforward network with 1024 intermediate hidden channels and attention with 512 hidden channels, 8 heads, as well as a static bias matrix of shape #$\#\operatorname{heads}\times N\times N$. All the metrics are recorded with a batch size of 1.
Figure 4: Efficiency ratio over “Pure FlashAttention”, which is calculated by Method EfficiencyPure FlashAttention $\frac{\text{Method Efficiency}}{\text{Pure FlashAttention Efficiency}}$.
@@ Results
According to Figure 3(a-b), when sequence length $N=16384$, FlashBias can significantly reduce GPU memory usage, which is 5$\times$ smaller than FlexAttention and FlashAttention with Bias during training and 10$\times$ smaller during inference. This memory benefit can also be theoretically justified by Theorem 3.2. As for training and inference running time, FlashBias presents 18.6% and 44% speedup compared to vanilla usage of FlashAttention and is also better than FlexAttention in long sequences.
@@ 4.2 Large Language Model
As stated in Example 3.4, we can adopt the exact decomposition for ALiBi bias, where $R$ is equal to 2 and the result of FlashBias is exactly equivalent to the original computation. Since attention masking has been well-optimized in FlashAttention dao2023flashattention and FlexAttention dong2024flex , we directly integrate them with our method to support the decoder-only Transformer. This combination also demonstrates that FlashBias is orthogonal to attention-masking speed-up techniques.
@@ Results
As mentioned above, since this task requires causal attention, we test FlashBias by integrating it with the attention mask speedup techniques in FlashAttention and FlexAttention. Results in Table 3 demonstrate that FlashBias still outperforms FlashAttention and FlexAttention in processing the ALiBi bias term. Specifically, in processing bias, FlashBias reduces over 50% (5.0$\to$2.3) time cost of FlashAttention and over 25% (3.4$\to$2.5) time cost of FlexAttention.
@@ 4.3 Vision Transformer
We test FlashBias on the image classification task based on Swin Transformer v2 liu2022swin . Specifically, we adopt the open-source SwinV2-B model [^1] Model and training configurations of SwinV2-B.. This model contains 24 layers with input resolution as $384\times 384$ and window size as $24\times 24$, thus, the sequence length of its WindowAttention is 576. WindowAttention in every layer contains a relative position bias with size #$\#\operatorname{heads}\times 576\times 576$, which is set as a learnable model parameter. We attempt to speed up the WindowAttention computation with FlashBias. All the efficiency metrics are evaluated under a batch size of 64.
@@
Since the biases are set as learnable parameters, we directly obtain them from a pretrained model and adopt the SVD decomposition to generate decomposed factor tensors. However, we observed that not all heads present low-rank bias matrices, as shown in Figure 5. Thus, we propose to split the heads into two subsets according to the biases’ ranks. For the subset with large ranks, we still adopt the vanilla FlashAttention with Bias, while adopting FlashBias for the low-rank subset.
@@
FlashBias (Ours) 87.126% 98.234% 0.190 5484
Results
In Table 4, we can find a serious performance drop in “Pure FlashAttention” (without bias, Acc@1 87$\to$9%), which verifies the essentiality of bias and also highlights the importance of speedup attention with bias. Further, compared to the official implementation, FlashBias can reduce 60% running time and 57% memory cost, which is valuable for real-time development. More importantly, FlashBias will only slightly affect the final performance, where the top-1 accuracy drops less than 0.02% and the fluctuation of the top-5 accuracy (0.002%) is within the standard deviation. In particular, even using well-established quantization techniques still will bring a 0.64% top-1 accuracy drop for 22% speedup according to the FasterTransformer document [^2] FasterTransformer released by NVIDIA.. This comparison demonstrates that the low-rank property is a principal basis for fast computation.
@@ 4.4 Scientific Deep Models
Table 5: Experiments of Transformer PDE solvers. The efficiency metrics are recorded under a batch size of 1 in the format of #Mem (GB) / #Time (s/100iters). Accuracy comparisons are in Appendix E.
@@ Results
After a detailed analysis of AlphaFold 3, we find that its efficiency bottleneck is triangle self-attention, which involves the bias matrix projected from intermediate pair representations, making it vary among different samples, layers and heads. To approximate this complex bias, we employ the neural decomposition version of FlashBias, whose inputs $\mathbf{x}_{\mathbf{q}},\mathbf{x}_{\mathbf{k}}$ are set as the combination of pair and single protein representation. We fine-tune neural basis functions for 2000 iterations on the PDB dataset. Since we only need to optimize the newly added parameters in $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ , this process only takes about 5 hours on a single A100 40GB GPU, then you can infer a new protein with FlashBias. For comparison, the full training process of AlphaFold 3 will take about 7 days on 128 A100 GPUs.
@@
Open-sourced Code 3.3724 0.9500 31.48 21.58
@@
Table 4.4 shows that, compared to the public code, FlashBias can reduce the running time and GPU memory usage by 30% and 50%, respectively. This speedup is very close to the efficiency upper bound: FlashAttention w/o Bias. Although FlashBias is based on the neural decomposition, it will not affect the final performance, whose metric fluctuation is within the standard deviation. Beyond inference, FlashBias can also be a promising tool for training speedup (see Appendix C).
@@ Appendix A Proofs for Theorems in the Main Text
Appendix B Speedup of Pangu-Weather bi2023accurate
@@
Since Pangu-Weather is based on the 3D window (with shape $2\times 6\times 12$), its bias for relative position encoding is slightly different from Swin Transformer. Especially, its bias is in the shape of #$\#\operatorname{num}\times\#\operatorname{heads}\times 144\times 144$ for each block, where #$\#\operatorname{num}$ represents the number of 3D windows. According to meteorological knowledge, different longitudes share the same bias.
@@ Results
Appendix C More Results in AlphaFold 3 abramson2024accurate
Potential in speeding up training
In the main text, we only evaluate FlashBias during the inference phase. Going further, as mentioned at the end of Section 3.2, if we directly replace the bias matrix in AlphaFold 3 abramson2024accurate with two decomposed factor functions, we can also accelerate the training process. As shown in Table 8, FlashBias can save around 9% running time and 30% GPU memory usage.
Table 8: Experiments of FlashBias in accelerating the training process of AlphaFold 3, where the default setting is to crop all the residue sequence into 384 tokens.
Method Running Time (s / 10 iters) GPU Memory (GB)
Open-sourced Code 153.272 23.552
FlashAttention with bias 150.058 22.740
FlashBias (Ours) 140.281 16.260

### v3

@@ Abstract
[…]
@@ 1 Introduction
In FlashAttention dao2022flashattention , researchers find that instead of computation FLOPS, the read-write (IO) overload of GPU high bandwidth memory (HBM) is the actual bottleneck of speed. As the bias matrix is usually dense, it is really hard to bypass the quadratic IO complexity, where the computation needs to read the whole bias term from HBM at least once, making its speedup intractable. In this paper, we notice that the IO challenge that we face here can be recast into the classical compressed sensing problem donoho2006compressed , whose basic assumption is that the “measurement” (corresponding to IO overload here) is expensive but the computation (corresponding to the fast on-chip computation) is cheap. This new perspective offers us valuable theoretical understandings. Specifically, for a dense matrix, the optimal “measurement” (IO overload) is highly related to the matrix’s rank candes2010power . This inspires us to dive into the rank of attention bias and surprisingly find that most of the widely used biases are inherently of low rank, eventually discovering one flexible way to enable fast computation of attention with bias.
@@
[…]
@@ 2.1 Attention with Bias
Additionally, large language models widely adopt rotary position embedding (RoPE) su2024roformer to introduce relative position into attention, which can be considered as a special multiplicative bias and is already computation efficient. Considering that RoPE is outside the additive bias scope, we defer its discussion to Appendix J and will showcase that FlashBias can also be extended to multiplicative bias.
@@ 3.1 Rethinking FlashAttention Computation
Given $\mathbf{q}\in\mathbb{R}^{N\times C},\mathbf{k},\mathbf{v}\in\mathbb{R}^{M\times C}$, bias $\mathbf{b}$ of rank $R$ and SRAM of size $S$ where $(C+R)\leq S\leq N(C+R)$ , there does not exist an algorithm to compute exact attention with bias through $o\big(\frac{NM(C^{2}+R^{2})}{S}\big)$ HBM access for all $S\in[(C+R),N(C+R)]$ . Here $o(\ast)$ represents the strict asymptotic upper bound.
@@ 3.2 FlashBias
where $[\ast|\ast]$ denotes the concatenation operation along the channel dimension. Notably, this design significantly reduces the storage cost for the attention bias from $\mathcal{O}(NM)$ to $\mathcal{O}\left((N+M)R\right)$. Although its design will require recalculating the bias weight, this computation is just a simple matrix multiplication of $\phi_{\mathbf{q}}(\mathbf{x}_{\mathbf{q}})\phi_{\mathbf{k}}(\mathbf{x}_{\mathbf{k}})^{\top}$, an operation that has been extremely optimized on modern GPUs.
@@
Some models such as Swin Transformer liu2021Swin and Pangu-Weather bi2023accurate adopt the learnable model parameters for relative position encoding. Specifically, each bias term in their model is an $N\times M$ matrix of model parameters. As this type of bias is fixed once the model has been well trained, it is convenient to conduct Singular Value Decomposition (SVD) klema1980singular for low-rank decomposition of these parameters. In practice, we precompute SVD once offline, incurring negligible runtime overhead. The resulting decomposed factor tensors can then be utilized to accelerate the subsequent inference process, thanks to their low-rank nature.
@@
Remark 3.6 (Understanding neural decomposition).
In Eq. (5), we define factor networks $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ as token-wise functions, instead of depending on the whole sequence. This is because elements in bias that reflect pair-wise relations are solely determined by corresponding token information, such as token position in relative position bias. Besides, rank $R$ essentially controls the dimension of dot-product space, thereby should be set based on “complexity” of prior knowledge that derives bias.
@@
Although SVD and neural decomposition will introduce approximation error, owing to the native low-rank property of attention bias in modern Transformers, it will not affect the final performance by setting $R$ as a reasonable value. For example, in some layers in Swin Transformer, $R=32$ can maintain over 99% energy of the original bias with shape $576\times 576$. More case studies are included in Appendix E and F.
@@
For FlashAttention with bias, its HBM access is $\Theta(\frac{NMC^{2}}{S}+NM)$ dao2023flashattention . Given a regular setting of Transformers and GPU ($C=64$, $S$ is 100KB) and supposing $R=64$, $N,M\gg C,R$, then HBM IO of FlashBias is $\approx 6\times$ smaller than FlashAttention.
@@ 4 Experiments
As summarized in Table 2, we extensively test FlashBias on Transformers for language modeling, computer vision and scientific problems, which can significantly speed up model efficiency at both training and test phases. This section will first present an overall efficiency comparison among different techniques and then integrate FlashBias to different domain-specific Transformers with various bias types. All experiments were performed in PyTorch 2.5.0 Paszke2019PyTorchAI and Triton 3.0.0 triton on a single A100 GPU.
@@ 4.1 Overall Comparison
To give a clear comparison among FlashBias, vanilla FlashAttention with Bias dao2023flashattention and the latest FlexAttention dong2024flex , we make a comprehensive efficiency evaluation based on a plain Transformer NIPS2017_3f5ee243 , which consists of 8 layers. Each Transformer layer involves a feedforward network with 1024 intermediate hidden channels and attention with 512 hidden channels, 8 heads, as well as a static bias matrix of shape $\operatorname{heads}\times N\times N$. All the metrics are recorded with a batch size of 1.
Figure 4: Efficiency ratio over “Pure FlashAttention”, which is calculated by $\frac{\text{Method Efficiency}}{\text{Pure FlashAttention Efficiency}}$.
@@ Results
[…]
@@
Figure 5: Speed of different implementations on attention with 128 hidden channels and 8 heads. $R$ is set as 8 in FlashBias. Dotted lines indicate out-of-memory. More attention speed comparisons are in Appendix B.
Implementation choices
As formalized in Eq. (3), for implementation, researchers can concatenate the decomposed factor functions with query and key tensors in the original attention along the channel dimension, which means FlashBias can seamlessly adopt previous efficient implementations for canonical attention, such as PyTorch SDPA [^1] PyTorch implementation of torch.nn.functional.scaled_dot_product_attention. In pursuit of a well-optimized operation, we also implement FlashBias based on the Triton backend triton . Figure 5 demonstrates that the Triton implementation is much more efficient in the forward pass and the PyTorch SDPA-based version is better in the training phase. Therefore, we adopt the Triton version for inference and PyTorch SDPA version for training throughout this paper. Besides, although vanilla PyTorch SDPA can accelerate the computation, it still suffers from the out-of-memory issue, which can be perfectly resolved by FlashBias.
@@ 4.2 Large Language Model
As stated in Example 3.4, we can adopt the exact decomposition for ALiBi bias, where $R$ is equal to 2 and the result of FlashBias is exactly equivalent to the original computation. Since attention masking has been well-optimized in FlashAttention dao2023flashattention and FlexAttention dong2024flex , we directly integrate them with our method to support the decoder-only Transformer. This combination also demonstrates that FlashBias is orthogonal to attention-masking speed-up techniques. Notably, FlashAttention has an ALiBi_slopes feature, which does not load bias matrix from HBM but computes the bias weights through just-in-time compilation. Since this is not a generic technique, we do not employ this feature in the main text. More discussion can be found in Appendix D.
@@ Results
[…]
@@ 4.3 Vision Transformer
We test FlashBias on the image classification task based on Swin Transformer v2 liu2022swin . Specifically, we adopt the open-source SwinV2-B model [^2] Model and training configurations of SwinV2-B.. This model contains 24 layers with input resolution as $384\times 384$ and window size as $24\times 24$, thus, the sequence length of its WindowAttention is 576. WindowAttention in every layer contains a relative position bias with size $\operatorname{heads}\times 576\times 576$, which is set as a learnable model parameter. We attempt to speed up the WindowAttention computation with FlashBias. All the efficiency metrics are evaluated under a batch size of 64.
@@
Since the biases are set as learnable parameters, we directly obtain them from a pretrained model and adopt the SVD decomposition to generate decomposed factor tensors. However, we observed that in one layer, not all heads present low-rank bias matrices, as shown in Figure 6. Thus, we only adopt FlashBias to layers where most of heads’ biases are of low rank, that is, the last 8 layers of SwinV2-B, while employing vanilla FlashAttention with Bias for the other layers.
@@ Results
FlashBias (Ours) 87.186% 98.220% 0.190 9429
More importantly, FlashBias will not affect the performance, where the top-5 accuracy drops less than 0.02% and the fluctuation of the top-1 accuracy (+0.042%) is within the standard deviation. In particular, even using well-established quantization techniques will still cause a 0.64% top-1 accuracy drop for 22% speedup according to the FasterTransformer document [^3] FasterTransformer released by NVIDIA.. This comparison demonstrates that the low-rank property is a principal basis for fast computation of attention.
@@ 4.4 Scientific Deep Models
Table 5: Experiments of Transformer PDE solvers. The efficiency metrics are recorded under a batch size of 1 in the format of #Mem (GB) / #Time (s/100iters). Accuracy comparisons are in Appendix G.
@@ Results
After a detailed analysis of AlphaFold 3, we find that its efficiency bottleneck is triangle self-attention, which involves the bias matrix projected from intermediate pair representations, making it vary among different samples, layers and heads. To approximate this complex bias, we employ the neural decomposition version of FlashBias, whose inputs $\mathbf{x}_{\mathbf{q}},\mathbf{x}_{\mathbf{k}}$ are set as the combination of pair and single protein representation. We fine-tune neural basis functions for 10,000 iterations on the PDB dataset. Since we only need to optimize the newly added parameters in $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ , this process only takes about 10 hours on a single A100 40GB GPU, then you can infer a new protein with FlashBias. For comparison, the full training process of AlphaFold 3 will take about 7 days on 128 A100 GPUs.
@@
Open-sourced Code 3.3724 0.9500 26.85 13.62
@@
Table 4.4 shows that, compared to the public code, FlashBias can reduce the running time by 32%. In addition, removing bias (w/o bias) will significantly improve the efficiency (26.85s vs. 8.27s) but will seriously damage performance. This observation highlights the essentiality of accelerating attention with bias. Despite FlashBias being based on neural decomposition, it will not affect the final performance, whose metric fluctuation is within the standard deviation. Beyond inference, FlashBias can also be a promising tool for training speedup (see Appendix E for details).
Neural decomposition visualization
To give a clear illustration of neural decomposition, we also plot the original pair representation bias and FlashBias approximated bias in Figure 7. Neural decomposition can give a relatively accurate estimation of the bias matrix, which performs well in capturing the “texture” of the original bias. In addition, it is also observed that neural decomposition is not completely perfect in reconstructing the diagonal weights. Despite this deficiency, FlashBias still maintains the original accuracy of AlphaFold 3. This may benefit from the dot-product self-attention and residual connection, which can give a robust and dominating weight for relation modeling.
Figure 7: Comparison between neural decomposed factor tensors’ multiplication and original bias. Biases of 7r6r (245 residues) and 7pzb (600 residues) in the first layer of Pairformer are plotted, which contains 4 heads. We also mark the rank value that can maintain 99% energy of original biases.
@@ 5 Conclusions
Acknowledgments and Disclosure of Funding
This work was supported by the National Natural Science Foundation of China (U2342217 and 62021002), the BNRist Innovation Fund (BNR2024RC01010), and the National Engineering Research Center for Big Data Software.
@@ Appendix A Proofs for Theorems in the Main Text
Appendix B Attention Speed Measurement
As a practical technique, in the main text, we primarily focus on the efficiency benefits to the entire Transformer brought by FlashBias. As a supplement, we conduct a detailed comparison solely on the efficiency of the attention mechanism’s computation. Specifically, we compared several advanced implementations of attention: FlashAttention dao2023flashattention , PyTorch SDPA [^5] PyTorch implementation of torch.nn.functional.scaled_dot_product_attention and xFormers xFormers2022 , as well as the vanilla PyTorch implementation. Since the official implementation of FlashAttention does not support backpropagation for bias matrix, we adopt an open-soured Triton extension for experiments [^6] Triton implementation of FlashAttention with bias..
Non-Causal Attention
[…]
Figure 8: Non-causal attention speed, where attention is with 4 heads and 32 hidden channels per head. $R$ is set as 8. The speed is measured with the batch size as 2.
Causal Attention
[…]
Figure 9: Causal attention speed, where attention is with 50 heads and 32 hidden channels per head. $R$ is set as 2. Dotted lines indicate out-of-memory. Speed is measured with the batch size as 1.
Appendix C Speedup of Pangu-Weather bi2023accurate
@@
Since Pangu-Weather is based on the 3D window (with shape $2\times 6\times 12$), its bias for relative position encoding is slightly different from Swin Transformer. Especially, its bias is in the shape of $\operatorname{num}\times\#\operatorname{heads}\times 144\times 144$ for each block, where $\operatorname{num}$ represents the number of 3D windows. According to meteorological knowledge, different longitudes share the same bias.
@@ Results
Appendix D More Results of ALiBi Bias press2022train
[…]
Table 8: Experiments on GPT-2 size model with ALiBi bias. Here we experiment with FlashAttention’s ALiBi_slopes feature and FlashBias with JIT speedup. The running time for 100 iterations at both training and inference phases when $N=2048$ are recorded.
Method Training Phase (s) Test Phase (s)
FlashAttention w/o bias 119.3 38.77
FlashAttention with ALiBi_slopes 119.8 38.98
FlashBias with decomposition in JIT 119.8 38.98
[…]
Appendix E More Results in AlphaFold 3 abramson2024accurate
Analysis of model efficiency
The theoretical complexities of triangle self-attention and triangle multiplication in AlphaFold 3 are both cubic w.r.t. the sequence length, while other components: data embedding, single attention with pair bias and feedforward layer, are of linear or quadratic complexity. Thus, according to theoretical complexity, these “triangle operations” are the bottleneck.
We further record the time cost of each component in AlphaFold 3 abramson2024accurate during the inference of PDB ID 7wux. As listed in Table 9, the triangle attention accounts for 53.3% of the total inference time.
Table 9: Running time of each component in AlphaFold 3 abramson2024accurate for inferring PDB ID 7wux.
Component Complexity w.r.t. Sequence Length Running Time (s) Ratio
Data Embedding Linear 1.91 7.1%
Triangle Self-attention Cubic 14.32 53.3%
Triangle Multiplication Cubic 9.97 37.1%
Single Attention with Pair Bias Quadratic 0.48 1.8%
FeedForward Linear 0.17 0.7%
Potential in speeding up training of AlphaFold 3 abramson2024accurate
In the main text, we only evaluate FlashBias during the inference phase. Going further, as mentioned at the end of Section 3.2, if we directly replace the bias matrix in AlphaFold 3 abramson2024accurate with two decomposed factor functions, we can also accelerate the training process. As shown in Table 10, FlashBias can save around 15.2% running time and 17.7% GPU memory usage compared to the open-sourced code.
Table 10: Experiments of FlashBias in accelerating the training process of AlphaFold 3 abramson2024accurate , where the default setting is to crop all the residue sequence within 384 tokens.
Method Running Time (s / 10 iters) Maximum GPU Memory (GB)
Open-sourced Code 165 23.572
FlashAttention with bias 153 23.572
FlashBias (Ours) 140 19.390